# references.bib

### v1

@@ 4.2 Experiment 2: Cross-dressed Characters
Lindabridis Female El castillo de Lindabridis 2842
Claridiana Female El castillo de Lindabridis 3332
Eugenia Female El José de las mujeres 3360
Rosaura Female La vida es sueño 3479
Semíramis Female La hija del aire II 4997
Table 6: Cross-dressing Characters
@@
We obtained a list of five cross-dressing characters from the database Role Call, [^7] https://rolecall.eu/crossdressed_characters/ listed in Table 6. All the characters are female characters cross-dressing as men: Lindabridis, Eugenia, Semíramis, Rosaura, and Claridiana. Each of these female characters subvert traditional roles by embodying manly traits such as intellectual aspirations, interest in war and hunting, physical strength, and resistance to marriage [araico2017lenguaje]. Information about in which scenes the characters are cross-dressing are taken both from the Role Call database and from looking at stage directions. Unfortunately there are many inconsistencies, and the cross-dressing is not always explicitly stated in the stage directions. In the cases where the stage direction indicates cross-dressing we use this information, otherwise we revert to the Role Call list.
@@
For the ecological validity of this experiment, we ensured that all cross-dressing characters were part of the validation set – that is, the gender classification model was not trained on any cross-dressing characters. As motivated above, we use the model that aggregates predictions at the scene level. We do not carry out any specific prediction for these character, but simply analyze the predictions of the model, as in Experiment 1, but now at the level of specific characters. We structure our analysis along two hypotheses.
Figure 3: Experiment 2: Confidence of the gender prediction model (aggregating at scene level) predicting the gender of vanilla and cross-dressing characters (green dot = correct prediction, red dot = incorrect prediction, green box = cross-dressing)
Hypothesis 1: Lower confidence.
@@
If the model picks up on gender differences portrayed by Calderón, the predictions for the cross-dressing should be less confident about their gender compared to other characters. Figure 3 shows the model’s overall confidence at the character level for all female characters from the four relevant dramas (cf. Table 6), with the five cross-dressers marked by green boxes. Compared to other female characters that speak the same number of words, the model predicted cross dressing characters with less confidence (mean = 0.78 compared to mean = 0.89). This makes sense, given that cross-dressing characters mix female-like and male-like speech. Additionally, the other female main characters were classified with slightly higher accuracy (4/5=0.80 compared to 19/23 =0.83); again, this is expected, given that we would expect a data-driven model to classify a cross-dressing character either way, depending on the extent of their cross-dressing. Indeed, the ’misclassified’ character is Claridiana, who cross-dresses in the whole play (see below).
@@
If the model does a good job in picking up gender cues in a fine-grained manner, it should be able to tell in which scene a character is cross-dressed. We tested this hypothesis, and Figure 4 shows the results for the four relevant dramas: The cross-dressing scenes (according the the role call database) are shown by a blue bar, and and the model predictions are indicated by the color of the scene number below (green: correct prediction, red: incorrect prediction). We find that, with a high proportion, cross dressing characters were predicted as male for the scenes in which they were cross-dressing — the model is able to pick up on instances in which these female characters are behaving in a stereotypically masculine fashion.
Figure 4: Experiment 2: Analysis of cross-dressing characters by scene. Cross-dressing scenes are indicated by blue bars. Model predictions are shown by the color of the act:scene number (red: incorrect prediction, green: correct prediction).
@@
The drama La hija del aire II focuses heavily on themes of governance and war as well as the mythical figure Semíramis, the queen of Asiria. Semíramis is described as being competent, a brave warlord, cruel and treacherous, but also evoking feminine sensuality. In the act three, after the death of her son, Semíramis disguises herself as her son in order to assume his place on the throne. In all of the scenes that Semíramis dresses as her son, the model predicts her as male, whereas through the rest of the play the model predicts her as female.
@@
In the story of El Castillo de Lindabridis, Lindabridis is a princess who travels the world in a flying castle looking for a husband. In this work, Claridiana is a female character who disguises as a male throughout the play in order to win the hand of Lindabridis. Claridiana engages in the trials of courage and combat required to marry Lindabridis. Claridiana is predicted by the model to be male with strong confidence in 14 out of 21 scenes in which she occurs. Lindabridis also disguises as a man in the third act of the play and in 6 out of 11 scenes of the third act she is predicted as male.
@@
Another clear case for the prediction model occurs in La vida es sueño, in which Rosaura disguises herself as a man in the first act of the play in an attempt to seek revenge on a man who took her honor. The cross-dressing is clearly picked up by the model because she is predicted as male in the majority of the first act, and solely as female in the rest of the acts.
@@
From these cases, we can see that not only is the the model is able to predict the gender of characters with high accuracy, it is also able to detect cases where female characters are cross-dressing. Therefore, we can see that the social context of the characters has a large influence on predictions of the model. The model is still able to predict many of these cross-dressing characters as women, but with lower confidence, indicating the complexity of the characters - they are women and embody traditional female traits, while using cross-dressing as a means to transcend their social role as women and accomplish tasks in the public social sphere, wield power, and take agency of their own stories[woods2013female].
@@ 5 Discussion and Conclusions
In this paper, we have taken a first step towards a ’scalable reading’ investigation of gender depiction on a large corpus of comedias written by Calderón de la Barca. [^8] https://anonymous.4open.science/r/calderon-gender-prediction-0A25/ We have proposed an analysis procedure where we train a gender classifier on the text (or parts of it) spoken by each character as an aggregation method. We then decompose the classifier’s predictions using an attribution method to critically assess the features that the classifier bases its predictions on. Since many attribution methods, including the integrated gradients we use here, are provably correct, this procedure can avoid the use of heuristic methods to identify keywords or topics, such as measures of association or keyness [keyness21] or topic modeling [10.1145/1553374.1553515]. Our method, nevertheless, keeps the domain expert at center stage and lets them determine the interpretation of the predictions and attributions presented.
@@
At the methodological level – what setups make this method work – we have investigated the best level at which to present the speech to the classifier. We find that presenting a person’s complete speech at one leads to robust results, but that classifying individual sentences and aggregating the results by probability does not only yield an improvement in overall accuracy, but also makes the predictions more fair for the minority gender (female). Despite their hunger for data, the models also perform reasonably well when asked to classify character speech from single scenes, allowing us to trace the plot in dramas involving cross-dressing characters.
@@
At the substantive level – what this method tells us about gender portrayal – we have shown that male and female characters broadly demonstrate adherence to gendered speech in a way that is classifiable at the sentence level. At the same time, our qualitative analysis on the example of Rosaura shows that, at the character level, Calderón’s gender representation can be non-monolithic and complex [arellano_ayuso_mujeres_2015] – and our computational tools provide us with a way of quantifying this intuition. Overall, in this way, our models produce findings that align with the existing hermeneutical studies, but can extend them to many of Calderón’s lesser known comedias. Concretely, we find that the main cues of gender identity that our models pick up regard modes of interaction (within-gender vs. cross-gender) as well as spheres of existence (the domestic and personal for female characters vs. the military and political for male characters).
@@
Another surprising but positive finding in our study was that several findings that looked wrong or at least counterintuitive at first glance turned out to be due to annotation errors in the corpus or the data bases we used, strengthening our belief in the accuracy of the method’s output and indicating its usefulness to uncover such errors in the annotation, an application known in computational linguistics as ’error mining’ [https://doi.org/10.1111/lnc3.12129]. [^9] We will correct the annotation errors we found and will submit these changes for publication with CalDraCor.
@@
More generally, our work indicates that Calderón’s dramas – at least in modernized and normalized versions – do not provide major obstacles to the use of current language model-based NLP, and thus lays the ground work for future use of scalable reading approaches in Calderónian studies.
@@
In terms of limitations, the study reported here was based on a gender classifier with a BERT architecture, which comes with a set of constraints. These include the use of a tokenizer (which may make some of the attributions difficult to understand) and the 512 token limit of the classifier. Most seriously, we limited ourselves to only the text spoken by each character, not taking into account dramatic structure or interpersonal relationships. Due to this simplification, the model is limited to lexical cues, and could not identify gender-based patterns like (hypothetically) ’men always speak first’. While our study shares this limitation with much work in stylometry [10.1093/llc/fqx023], it is possible to present such structural information to classifiers as well [kim-klinger-2019-frowning], in particular in more recent models with longer context windows [NEURIPS2020_c8512d14]. We leave such endeavors to future work.
@@
We conclude by sketching two more avenues for future research. The first one is to move beyond gender to other character traits such as socio-econnomic status. Honor is a key theme in the works of Calderón, and it intimately tied to socio-economic status and hierarchy [greer2017play][ibarreche2020honor]. Particularly, noble characters spend many plays discussing the losing and gaining of honor. Characters of high socio-economic status are more likely to be well educated and thus speak differently to lower class character or servants [woods2013female]. Siglo de Oro authors also utilized character archetypes which tie to different socio-economic statuses such as the gracioso - a servent character who serves as comic relief. These are frequently labeled in the cast sheet, however not all instances of this character type are annotated in our corpus. Therefore, it may be possible to utilize the same classification method to identify socio-economic status and enrich the annotations in CalDracor. Identiyfing archetypes may require the use of additional processing tools, such clustering. This information can then be used to identify structural patters, in Calderon’s work.
@@
A second direction could be the use of our approach to find structure in the large number of works by Calderón, looking for example at differences in gender portrayal between his tragedies and his comedies, or conversely using gender depiction as evidence to argue for or against certain genre delineations [lehmann23:_clasif_traged_comed_comed_nuevas_calder_barca]. Another possible question to ask is whether Calderón’s gender portrayal remained constant over the long period of his productive career, or whether it changed.
@@ 6.1 Model training details
In order to use the model for our gender classification task, we add a binary classification head on top. We use the character lines from CalDraCor together with the ground truth gender labels as data to tune this classification head. We used the Adam optimizer, a standard gradient descent based optimizer [zhang2018improved], and considered a range of reasonable hyperparameter values recommended in the literature. The best values for each models are listed in Table 7 This resulted in a learning rate of 1e-05 for each model. We implemented early stopping to prevent overfitting of the training data.
@@ 6.2 Indicative tokens: English translation
(’comarad’, 0.5076) 2 (’##acas’, -0.3066) 1
(’accomodations, 0.4962) 3 (’discourses, speeches’, -0.2966) 1

### v2

@@ 4.2 Experiment 2: Cross-Dressed Characters
Rosaura Female La vida es sueño 3,479
Eugenia Female El José de las mujeres 3,360
Lindabridis Female El castillo de Lindabridis 2,842
Semíramis Female La hija del aire II 4,997
Claridiana Female El castillo de Lindabridis 3,332
Table 6: Cross-dressing Characters.
@@
We obtained a list of five cross-dressing characters in four of Calderón’s plays, listed in Table 6, from the Role Call database. [^7] See: https://rolecall.eu/crossdressed_characters/. Since we are not aware of other sources for instances of cross-dressing in Calderón’s work, we treat this list as ground truth, although the possibility remains that the list is not exhaustive. All five characters are female characters cross-dressing as men: Rosaura, Eugenia, Semíramis, Claridiana, and Lindabridis. Each of these female characters subvert traditional roles by embodying manly traits such as intellectual aspirations, interest in war and hunting, physical strength, and resistance to marriage ([araico_2017]). Information about in which scenes the characters are cross-dressing are taken both from the Role Call database and from looking at stage directions. Unfortunately, there are many inconsistencies, and the cross-dressing is not always explicitly stated in the stage directions. In the cases where the stage direction indicates cross-dressing, we use this information, otherwise we revert to the Role Call list.
@@
For the ecological validity of this experiment, we ensure that all cross-dressing characters are part of the validation set – that is, the gender classification model is not trained on any cross-dressing characters. As motivated above, we use our trained model that aggregates predictions at the scene level. We do not carry out any specific prediction for these characters, but simply analyze the predictions of the model, as in Experiment 1, but now at the level of specific characters. We structure our analysis along two hypotheses.
Figure 3: Experiment 2: Probability that the character is predicted as female (prediction made at our first input level, all lines) for all female characters speaking more than 2000 words (green dot = predicted as female, red dot = predicted as male, green box = cross-dressing). Note: The highlighted instance of Semíramis is from la Hija del Aire II.
Hypothesis 1: Lower Confidence for Cross-Dressers.
@@
If the model picks up on gender differences used strategically by Calderón, the predictions for the cross-dressing characters should be less confident about their gender than for other characters. Figure 3 shows the model’s overall confidence at the character level for all female characters from the four relevant dramas (Table 6), with the five cross-dressers marked by green boxes. Compared to other female characters that speak the same number of words, the model predicted cross-dressing characters to be female with less confidence (mean = 0.78 compared to mean = 0.89). Only four female characters are predicted with comparably low confidence (Clariana, Leonor, Celia, and Aurora), all speaking comparatively little. This pattern makes sense, given that cross-dressing characters mix female-like and male-like speech, providing conflicting evidence to the model.
Additionally, the model struggles with the overall classification of the cross-dressed characters more than with non-cross dressing ones (the accuracy is 4/5=0.80 compared to 19/23 =0.83). Again, this is expected, given that we would expect a data-driven model to classify a cross-dressing character either way, depending on the extent of their cross-dressing. Indeed, the ‘misclassified’ character is Claridiana, who cross-dresses in the whole play (see below).
@@
If the model does a good job of picking up gender cues in a fine-grained manner, it should be able to tell in which scene a character is cross-dressed. We test this hypothesis on the four relevant dramas, and Figure 4 shows the results: The cross-dressing scenes (according to the Role Call database, and in some cases identified by us in the few instances where stage directions indicate cross-dressing) are shown by a blue shading, and the model predictions are indicated on the y axis: the probability that a character was predicted to be female. Along the x axis are all the scenes in which a character speaks (ex. 1:2 on the axis represents Act 1, Scene 2). We find that, with a proportion higher than chance, cross-dressing characters were predicted as male for the scenes in which they were cross-dressing – the model is able to pick up on instances in which these female characters are behaving in a stereotypically masculine fashion.
Figure 4: Experiment 2: Analysis of cross-dressing characters by scene. Cross-dressing scenes are indicated by blue bars. y axis represents the probability that the character will be predicted as female.
@@
The drama La hija del aire II focuses heavily on themes of governance and war as well as the mythical figure Semíramis, the queen of Asiria. Semíramis is described as being competent, a brave warlord, cruel and treacherous, but also evoking feminine sensuality. In the act three, after the death of her son, Semíramis disguises herself as her son in order to assume his place on the throne. In all of the scenes that Semíramis dresses as her son, the model predicts her as male, whereas through the rest of the play the model mostly predicts her as female.
@@
In the story of El Castillo de Lindabridis, Lindabridis is a princess who travels the world in a flying castle looking for a husband. In this work, Claridiana is a female character who disguises as a male throughout the play in order to win the hand of Lindabridis. Claridiana engages in the trials of courage and combat required to marry Lindabridis. Claridiana is predicted by the model to be male in 14 out of 21 scenes in which she occurs. Lindabridis also disguises herself as a man in the third act of the play and in 6 out of 11 scenes of the third act she is predicted as male.
@@
Finally, the model does a good job in La vida es sueño, in which Rosaura disguises herself as a man in the first act of the play in an attempt to seek revenge on a man who took her honor. The cross-dressing is clearly picked up by the model because she is predicted as male in the majority of the first act (4/5 scenes) and mostly as female in the rest of the acts (8/11 scenes).
From these cases, we can see that not only is the the model able to predict the gender of characters with high accuracy, it is also able to detect specific scenes where female characters are cross-dressing. Therefore, we can see that the social context of the characters has a large influence on the predictions of the model. Still, a more fine-grained analysis of the individual plays would help contextualize the model’s mispredictions at the scene level.
@@
In any case, the model is still able to predict many of these cross-dressing characters as women, but with lower confidence, indicating the complexity of the characters – they are women and embody traditional female traits, while using cross-dressing as a means to transcend their social role as women and accomplish tasks in the public social sphere, wield power, and take agency of their own stories ([woods_2013]).
@@ 5 Discussion and Conclusions
In this paper, we have taken a first step towards a ‘scalable reading’ investigation of gender depiction on a large corpus of comedias written by Calderón de la Barca. We make our implementation publicly available. [^8] Software can be found here: https://github.com/allisonakeith/calderon-gender-prediction (DOI: https://doi.org/10.5281/zenodo.15639883). We have proposed an analysis procedure where we train a gender classifier on the text (or parts of it) spoken by each character as an aggregation method. We then decompose the classifier’s predictions using an attribution method to critically assess the features that the classifier bases its predictions on. Since many attribution methods, including the integrated gradients we use here, are provably correct, this procedure can avoid the use of heuristic methods to identify keywords or topics, such as measures of association or keyness ([schroeter_2021, vsecla_2024]) or topic modeling ([wallach_2009]). Our method, nevertheless, keeps the domain expert at center stage and lets them determine the interpretation of the predictions and attributions presented.
@@
The contributions for computational linguistics lie at the methodological level, in better understanding what properties of the setup are important for the success of the method. We investigate the best level at which to present the speech to the classifier and find that presenting a person’s complete speech at once leads to robust results, but that classifying individual sentences and aggregating the results by probability does not only yield an improvement in overall accuracy, but also makes the predictions more accurate for the smaller gender category (female). Despite their hunger for data, the models also perform reasonably well when asked to classify character speech from single scenes, allowing us to trace the plot in dramas involving cross-dressing characters.
@@
The contribution for literary studies lie at the substantive level, using the method to gain a better understanding of gender portrayal in Calderón’s comedias. We have shown that male and female characters broadly demonstrate adherence to gendered speech in a way that is classifiable at the sentence level. Our findings bolster prior observations by [vsecla_2024], finding similar thematic differences between male and female characters (female characters making reference to their internal worlds, for example). At the same time, our qualitative analysis on the example of Rosaura shows that, at the character level, Calderón’s gender representation can be non-monolithic and complex ([arellano_2015]) – and our computational tools provide us with a way of quantifying this intuition. Overall, in this way, our models produce findings that align with the existing hermeneutical studies, but can extend them to many of Calderón’s lesser known comedias. Concretely, we find that the main cues of gender identity that our models pick up regard modes of interaction (within-gender vs. cross-gender) as well as spheres of existence (the domestic and personal for female characters vs. the military and political for male characters).
@@
Another surprising but positive finding in our study was that several findings that looked wrong or at least counterintuitive at first glance turned out to be due to annotation errors in the corpus or the data bases we used, strengthening our belief in the accuracy of the method’s output and indicating its usefulness to uncover such errors in the annotation, an application known in computational linguistics as “error mining” ([dickinson_2015]). [^9] We will correct the annotation errors we found and will submit these changes for publication with CalDraCor.
@@
More generally, our work indicates that Calderón’s dramas – at least in modernized and normalized versions – do not provide major obstacles to the use of current language model-based NLP, and thus lays the ground work for future use of scalable reading approaches in Calderonian studies.
@@
In terms of limitations, the study reported here was based on a gender classifier with a BERT architecture, which comes with a set of constraints. These include the use of a model-specific tokenizer which cannot be easily adapted to the specific corpus and which may make some of the attributions difficult to understand, as well as the 512-token limit of the classifier. Most seriously, we limited ourselves to only the text spoken by each character, not taking into account dramatic structure or interpersonal relationships. Due to this simplification, the model is limited to lexical cues, and could not identify gender-based patterns like (hypothetically) ‘men always speak first’. While our study shares this limitation with much work in stylometry ([evert_2017]), it is possible to present such structural information to classifiers as well ([kim_2019]), in particular in more recent models with longer context windows ([zaheer_2020]). We leave such endeavors to future work.
@@
We conclude by sketching three more avenues for future research. The first one is to move beyond gender to other character traits such as socio-economic status. Honor is a key theme in the works of Calderón, and it is intimately tied to socio-economic status and hierarchy ([greer_2017, ibarreche_2020]). Particularly, noble characters spend many plays discussing the losing and gaining of honor. Characters of high socio-economic status are more likely to be well educated and thus speak differently to lower class character or servants ([woods_2013]). Siglo de oro authors also utilize character archetypes which are tied to different socio-economic statuses such as the gracioso – a servant character who serves as comic relief. These are frequently labeled in the cast sheet, however not all instances of this character type are annotated in our corpus. Therefore, it may be possible to utilize similar classification methods to identify socio-economic status and enrich the annotations in CalDraCor. Identifying archetypes may require the use of additional processing tools, such as clustering. This information can then be used to identify structural patters, in Calderón’s work.
@@
A second direction could be the use of our approach to find structure in the large number of works by Calderón, looking for example at differences in gender portrayal between his tragedies and his comedies, or conversely using gender depiction as evidence to argue for or against certain genre delineations ([lehmann_2023]). Another possible question to ask is whether Calderón’s gender portrayal remained constant over the long period of his productive career, or whether it changed.
Finally, future work could apply the methods developed here to other dramas by siglo de oro authors such as Lope de Vega. This would permit us to understand the extent to which gender portrayal we observe in Calderón’s comedias is a general feature of the period or author specific. This would require overcoming practical hurdles, though, since other authors’ work are generally not as easily accessible and consistently represented as Calderón’s in CalDraCor.
Acknowledgements.
This work was conducted as part of the ’Identifying Regularities in the works of Pedro Calderón de la Barca’ project (508056339) funded by the DFG Priority Programme / Schwerpunktprogramm ’Computational Literary Studies’ (SPP 2207).
@@ 6.1 Model training details
In order to use the model for our gender classification task, we add a binary classification head on top. We use the character lines from CalDraCor together with the ground truth gender labels as data to tune this classification head. We used the Adam optimizer, a standard gradient descent based optimizer [zhang_2018], and considered a range of reasonable hyperparameter values recommended in the literature. The best values for each models are listed in Table 7 This resulted in a learning rate of 1e-05 for each model. We implemented early stopping to prevent overfitting of the training data.
@@ 6.2 Indicative tokens: English translation
(’comrade’, 0.5076) 2 (’##acas’, -0.3066) 1
(’camp, accommodations, 0.4962) 3 (’discourses, speeches’, -0.2966) 1